hey once again i tried this problem a few times and the answer doesnt match what the teacher gave me as the answer.....
factor perfect-square trinomials and differences of squares.....
X^2Y-4Y+X^2-4
=y(x^2-4)+(x^2-4)
=(x^2-4)(y+1)
=(x-2)(x+2)(y+1)
Cheers

Answer by stanbon(75887)

x^2y-4y+x^2-4
=y(x^2-4)+(x^2-4)
=(x^2-4)(y+1)
=(x-2)(x+2)(y+1)
Cheers,
Stan H.